OpenStudy (anonymous):

the area of a circle with radius r is given by the formula A=πr^2. if a circle has area 64π square inches, what is the radius?

OpenStudy (anonymous):

You have been given A = 64*pi and also that A = pi*r^2 . plug in value of A in the above equation , that gives you 64*pi = pi*r^2 divide both sides of equation by pi, which gives you 64 = r^2 therefore , \[r=\sqrt{64}\] inches